OpenStudy (anonymous):

The table shows numbers with the ratio x : y. Find the missing number. x 1 15 225 y 4 ? 900 A. 30 B. 60 C. 80 D. 90

OpenStudy (anonymous):



OpenStudy (anonymous):

the ratio is the same for each entry. so 4/1 = 900/225 = y/15

OpenStudy (anonymous):

solve y/15 =4